OpenStudy (anonymous):

The profits in hundreds of dollars, P(c), that a company can make from a product is modeled by a function of the price, c, they charge for the product: P(c) = –20c2 + 320c + 5,120. What is the maximum profit the company can make from the product?

OpenStudy (anonymous):

divide by 20 first to get -c^2+16c+256 max = -b/2a =-16/2=8 the minus showing a maximum P(c)=P(8)= -64+144+256=320 so coords.of max (8,320) multiply by 20 to compensate for original division (160,6400)